two air planes leave at the same time traveling in the same direction. At six hours they are 1530 miles apart. 1st plane is 4 times faster than the 2nd plane. I am to find how fast each plane is going. "??????"
I can not get anywhere with this problem. plane A is 4X and plane B is X. I do not know distance except for they are 1530 apart after 6 hours. I am stuck.
No help from my math book and this problem is due by tomorrow. Help!!!

Answer by stanbon(75887)

two air planes leave at the same time traveling in the same direction. At six hours they are 1530 miles apart. 1st plane is 4 times faster than the 2nd plane. I am to find how fast each plane is going.
----------------------
Slower plane DATA:
rate = x mph ; time = 6 hrs ; distance = r*t = 6x miles
-----
Faster plane DATA:
rate = 4x mph ; time = 6 hrs ; distance = r*t = 24x miles
-----
Equation:
faster distance - slower distance = 1530 miles
24x - 6x = 1530
---
18x = 1530
---
x = 85 mph (slower plane speed)
4x = 340 mph (faster plane speed)
================
Cheers,
Stan H.